OpenStudy (anonymous):

Fan and medal a. A bike tire has a pressure of 3.2 atm at 300 K. What would the pressure be if the temperature were lowered to 250 K? b. A balloon has a volume of 18.6 L at sea level (1 atm). As you walk down a valley, the pressure changes to 0.854 atm. How does this affect the volume of the balloon? c. The internal conditions for a half drank water bottle are measured to be 18OC and 1.2 atm. How would the pressure change if the temperature were lowered to -3OC.

OpenStudy (matt101):

I'll show you how to do the first one, and that should make the others pretty easy! For a) - use Amonton's Law: \[\frac{ P _{1} }{ T _{1} }=\frac{ P _{2} }{ T _{2} }\]\[\frac{ 3.2 }{ 300 }=\frac{ P _{2} }{ 250 }\]\[P_{2}=2.7\] For b) - use Boyle's Law: \[P _{1}V_{1}=P_{2}V_{2}\] For c) - use Amonton's Law again, but make sure you convert temperature to Kelvin (add 273).